# Atomic forces from correlation energy functionals based on the adiabatic- connection fluctuation-dissipation theorem

### v1

@@ I Introduction
The analytical forces in the RPA have been considered in a few works. In Refs. [24, 25, 26, 27, 28, 29], the analytical non-self-consistent RPA forces (i.e. the analytical gradient of the RPA energy evaluated with orbitals and eigenvalues from a semi-local approximation) was implemented in the context of Gaussian basis sets, and tested on small molecules. Recently, Thierbach et al. [30, 31] used Gaussian basis sets to evaluate RPA forces at self-consistency, solving the optimized effective potential (OEP) equation [32, 33, 34, 35, 36, 37, 38, 39]. While Gaussian functions are efficient for describing finite-size systems such as molecules, the use of plane waves and pseudopotentials is often more convenient in solids. To date, only one work has reported non-self-consistent RPA forces with plane waves in the projector-augmented-wave (PAW) framework [40].
In this work, we derive analytical expressions for the self-consistent and non-self-consistent evaluation of RPA forces, in the presence of nonlocal external potentials, such as those arising from norm-conserving pseudopotentials [41]. The force expressions are implemented in the context of plane-wave basis sets within the ACFDT package [42, 36, 12, 43, 14, 18, 38] of the Quantum ESPRESSO (QE) code [44, 45], which exploits an eigenvalue decomposition of the linear density response function for evaluating the RPA energy and potential. We demonstrate that our implementation achieves a high numerical quality, enabling the exploitation of RPA forces in a wide range of applications. We also determine forces within RPAx using finite-difference total energy calculations. The accuracy of the different methods is assessed by comparing results for structural and vibrational properties of molecules and solids to published results obtained with coupled-cluster single, double, and perturbative triple excitations (CCSD(T)), Diffusion Monte Carlo (DMC), and experiments.
The paper is organized as follows. In Sec. II, we recall the ACFDT total energy expression. In Sec. III, we derive the RPA atomic forces in detail and discuss their implementation. In Sec. IV, we demonstrate the numerical accuracy of the calculated forces. In Sec. V, we present the performance of the RPA and RPAx methods in predicting structural and vibrational properties of a set of molecules (H2, LiH, FH, N2, H2O, NH3, and CH4) and solids (C, Si, and Ge). Finally, we present our conclusions in Sec. VI.
@@ II ACFDT formalism
Within the Born-Oppenheimer (BO) approximation, the potential energy surface is given by the sum of the following contributions: the internuclear interaction energy, $E_{\text{NN}}$, the electronic kinetic energy, $T_{e}$, the electron-nuclei (external) interaction energy, $E_{\text{ext}}$, and the electron-electron interaction energy, $W$. Within the KS-DFT framework, the sum of $W$ and $T_{e}$ is re-expressed as a sum of the kinetic energy of independent KS electrons, $T^{\rm KS}_{e}$, the Hartree energy, $E_{\text{H}}$, the exchange energy, $E_{\text{x}}$, and the correlation energy, $E_{\text{c}}$. Typically, as in the LDA and GGA, both $E_{\text{x}}$ and $E_{\text{c}}$ are approximated.
Approximations such as the RPA and RPAx have an exact description of the exchange energy, while the correlation energy is approximated starting from the formally exact ACFDT formula [6, 7]
$E_{\text{c}}^{\text{ACFDT}}=\int\limits_{0}^{1}d\lambda\int_{0}^{\infty}\frac{du}{2\pi}\,\text{Tr}\bigl\{$ $v\bigl[\chi_{\lambda}(iu)-\chi_{\text{0}}(iu)\bigr]\bigr\}.$ (1)
Here, $v$ is the bare Coulomb repulsion and $\lambda$ is a parameter connecting the fictitious KS system of independent particles ($\lambda\!=\!0$) to the physical system ($\lambda\!=\!1$), while keeping the charge density fixed at $\lambda=1$.
The key ingredient in the ACFDT formula is the linear density response function, evaluated for the $\lambda$-scaled interacting system and for the KS system. The latter is defined by the one-particle KS potential
@@
The two response functions, $\chi_{\lambda}$ and $\chi_{\text{0}}$, are connected through the Dyson equation of time-dependent density functional theory [46]
@@
The Hartree and exact-exchange kernels are both linear in $\lambda$ ($f^{\lambda}_{\text{x}}=\lambda f_{\text{x}}$). It is, therefore, possible to integrate over $\lambda$ in Eq. (1), analytically. Combining Eqs. (1) and (3) gives the following expression for the correlation energy in RPAx (or RPA, by setting $f_{\text{x}}=0$)
@@
Note here that $f_{\text{x}}$ is frequency-dependent [48]. In Refs. [12, 14], the use of alternative partial resummations of the RPAx was motivated. Using the same $f_{\text{x}}$-kernel, we can define RPAx(1) from
@@
which corresponds to a truncation of the RPAx(1) series as it omits exchange in the screened interaction. Once the $f_{\text{x}}$-kernel is calculated, both variants are straightforwardly obtained. They usually give very similar results to RPAx but, in certain critical cases, the alternative resummations are preferred.
The response functions involve summations over excitation functions, i.e., products of occupied and unoccupied KS states. In Ref. [42], it was shown that the use of unoccupied states can be completely avoided for the RPA correlation energy by solving the following eigenvalue problem
@@
Determining the RPA and RPAx correlation energy requires a starting point. Such non-self-consistent calculations always introduce a bias, which can only be completely suppressed by identifying the optimal KS starting point. This is accomplished through the optimized effective potential (OEP) approach [32, 33, 34, 35, 36, 37, 38, 39], which allows orbital-dependent functionals to work within the standard Kohn-Sham scheme of DFT [51]. It applies not only to RPA but also to any method that includes exact exchange, such as hybrid functionals. For a given xc functional, the OEP equation reads
@@
Both the left and right-hand sides of Eq. (9) can be written in terms of density responses. This feature was exploited in Ref. [36] to solve the OEP equation, combining DFPT with an iterative algorithm for extracting the xc potential. The RPA correlation term contributing to the left-hand side of Eq. (9) can be evaluated as
@@
The variation of each eigenvalue with respect to $v_{\text{eff}}$ can be split further into different terms, which can all be evaluated using DFPT [36, 38].
Following the eigenvalue problem in Eq. (7), the total energy in RPA/RPAx, and the potential of RPA have been implemented within the ACFDT code [42, 36, 12, 43, 14, 18, 38] of the QE distribution [44, 45]. Building on these capabilities, we have, in this work, implemented the forces as described in the next section.
@@ III Atomic forces in RPA
Having established a way to calculate the total energy, we can now look at its first-order derivatives with respect to the atomic positions. This yields a set of forces, which are important for studying the stability and excitations of the lattice or a molecular structure. The force, $\mathbf{F}_{I}$, exerted on a given atom, $I$, corresponds to the derivative of the BO potential energy surface with respect to the atomic position, $\mathbf{R}_{I}$. Thus, from Sec. II, we have
$\mathbf{F}_{I}=-\frac{\partial E_{\text{NN}}}{\partial\mathbf{R}_{I}}-\frac{\partial T_{e}}{\partial\mathbf{R}_{I}}-\frac{\partial E_{\text{ext}}}{\partial\mathbf{R}_{I}}-\frac{\partial E_{\text{H}}}{\partial\mathbf{R}_{I}}-\frac{\partial E_{\text{xc}}}{\partial\mathbf{R}_{I}}.$ (11)
@@
is the external (nuclear) potential. The first term in Eq. (12) corresponds to the Ewald contribution, which is linked to the internuclear repulsion [52]. The second term describes the interaction between the nuclei and the electronic density. The third and fourth terms are the ones that depend directly on the xc functional. In the next two subsections, we will adapt Eq. (12) to RPA at self-consistency and to non-self-consistent RPA, starting from a local or semi-local approximation.
@@ III.1 Forces at self-consistency
As mentioned previously, in order to avoid starting-point biases, a self-consistent (scf) solution should ideally be considered. Furthermore, self-consistency simplifies the expression for the force given in Eq. (12), thanks to the validity of the Hellmann-Feynman theorem (HFT) [53, 54]. At self-consistency, the xc potential is equal to the functional derivative of the xc energy with respect to the charge density. By applying the chain rule on the fourth term in Eq. (12), and using the fact that, at self-consistency, the density and xc potential are identical in both the third and fourth term, it is easy to see that these two contributions cancel. We can thus write
@@
This is the textbook Hellmann-Feynman force formula, used within the PW code of QE to compute atomic forces at the LDA and GGA levels of theory.
@@
The effective KS potential $v_{\text{eff}}(\mathbf{r}^{\prime},\mathbf{r}^{\prime\prime})$ is now allowed to have a (static) nonlocal component coming from Eq. (15). Focusing on the RPA functional, the fourth term of Eq. (12) becomes
@@
The collection of eigenvalues $\{a_{\alpha}\}$ and their variations are obtained directly from the calculation of the RPA correlation energy and potential. The detailed expression for $\delta a_{\alpha}(iu)/\delta v_{\text{eff}}$ when $v_{\text{eff}}$ is local is given in Refs. [36, 38]. For a nonlocal $v_{\text{eff}}$, as in the present case, off-diagonal components also need to be calculated. Fortunately, this does not cause any additional complication. As a result, implementing the calculation of $\partial a_{\alpha}(iu)/\partial\mathbf{R}_{I}$ in the existing scf implementation is relatively straightforward. The computation of the extra scf-RPA correlation force term is also convenient as it can be done together with the computation of the RPA correlation potential. Furthermore, since only the pseudopotential is involved in Eq. (18), the density and eigenvalue variations correspond to bare responses, which facilitate their evaluation.
@@ III.2 Non-self-consistent forces
The force expression in Eq. (12) is general and valid for any KS starting point. For ease of discussion, we will hereafter focus on a particular starting point, the PBE approximation [57, 58]. In this case, the second term in Eq. (12) is evaluated with the PBE density, and the third force term is evaluated with the PBE xc potential. The derivative of $n^{\text{PBE}}$ with respect to $\mathbf{R}_{I}$ is determined from the first-order variation of the occupied PBE orbitals, which can be computed from the self-consistent Sternheimer equation,
@@
where $\hat{H}_{\rm KS}$ stands for the PBE KS Hamiltonian. For a given occupied band $\nu$, the quantity $\varepsilon_{\nu}$ corresponds to its KS eigenvalue and $\Delta\phi_{\nu}$ is its first-order response to the applied perturbation, $\Delta\mathbf{R}_{I}$. The quantity $\Delta\hat{V}$ stands for the response of the KS potential to the same perturbation. We note that, since self-consistency is required, the only feasible starting points are those from LDA and GGA functionals, for which the DFPT formalism is implemented within QE [50].
@@
Figure 1: Change in the total force for the fluorane molecule according to a) HF, EXX, and the uncorrected EXX (${\bf F}^{\rm scf\text{-}EXX}_{\rm x}$), b) RPA, RPA[PBE], and the uncorrected RPA (${\bf F}^{\rm scf\text{-}RPA}_{\rm xc}$).
@@ IV Implementation and testing
In order to judge whether the quality of the forces are sufficient for relaxing geometries using, for example, the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm [62, 63, 64, 65], we have considered the HF molecule. The hydrogen atom has no core electron and local PPs are found to converge well. On the other hand, fluoride has two core electrons and, as such, benefit from using nonlocal PPs.
@@
$\text{F}_{\text{tot}}=\sqrt{\sum_{I=1}^{M}\sum_{\beta=1}^{3}(F_{I}^{\beta})^{2}},$ (23)
@@
where $M$ is the total number of atoms and $\beta$ spans the three spatial Cartesian coordinates. The total force takes into account the individual forces exerted on both the fluoride and hydrogen atoms, and should tend toward zero as the optimal geometry is approached. The change in the total force, calculated at different values of the F-H bond distance, is given in Fig. 1, for different approximations. Panel a) focuses on the methods based on HF. The notation ’HF’ refers to a self-consistent calculation using the nonlocal HF potential, as implemented within the PW program of QE. The ’HF[PBE]’ method corresponds to its non-self-consistent version that we have implemented within the ACFDT package, and ’EXX’ refers to the self-consistent local version of HF that solves the OEP equation. Finally, ’EXX $\{-\mathbf{F}_{\text{x}}^{\text{scf-EXX}}\}$’ corresponds to an EXX calculation that omits the extra exchange force contribution derived in Ref. [41] (exchange part of Eq. (18)). For all these calculations, we used a kinetic energy cutoff of 100 Ry and a simulation cell of size 20 bohrs. For the HF and EXX calculations, a tight self-consistency threshold was set in order to ensure a good convergence of the results. The same was done for obtaining the self-consistent solution to the Sternheimer equation, necessary for computing the forces at the HF[PBE] level (see Eq. (20)).
@@
The FH molecule is found to have an equilibrium bond distance of 0.8924 Å using HF[PBE], 0.8948 Å for EXX, and 0.8952 Å for HF. The minimum in $\text{F}_{\text{tot}}$ is found, for all three methods, at the same position as the minimum in total energy, as expected. However, when we omit the extra exchange force term due to the nonlocal PP, the optimal bond length is about 0.03 Å larger than expected, with an overall error in the total force of about 0.085 Ry/a0. When inspecting the individual forces exerted on each atom along a given Cartesian direction, the weight of the extra PP term is found, as expected, larger for the fluoride atom than the hydrogen, 0.11 Ry/a0 against $\times1$0-4 Ry/a0, respectively.
@@
Panel b) in Fig. 1 focuses on RPA using the same computational setup as for the EXX calculations. For the evaluation of the correlation terms, the frequency integration was performed on a grid of 8 imaginary frequency points, and 10 eigenvalues per valence electron ($N_{\rm eig}$=80) of the KS response function were determined (see Eq. (7)). The bond distance is found at 0.9223 Å with RPA[PBE] and 0.9224 Å with RPA. In both cases, the minimum in the total force matches the minimum associated with the total energy, as expected. When removing the extra xc force term (RPA $\{-\mathbf{F}_{\text{xc}}^{\text{scf-RPA}}\}$), the bias in forces leads to a shorter F-H bond. The error in the evaluation is about 0.01 Å, which corresponds to a shift of 0.027 Ry/a0 in the total force. Again, the extra xc force term is found to be the largest for F, about 0.04 Ry/a0, while it is negligible for H.
@@
Given the good agreement found at equilibrium geometry using either the force or the total energy, we then investigated how well our calculated forces are exploited by the BFGS algorithm to relax the geometry of fluorane. For this test, the starting F-H bond distance was set to 1.2 Å. During the relaxation procedure, we fixed the position of the hydrogen atom and only let the position of F to vary. The change over the relaxation steps in the total energy, the total force, and the F-H bond distance, are presented in panels a), b), and c) of Fig. 2, respectively.
@@
Figure 2: Evolution of a) the total energy, b) the total force, and c) the bond distance during the BFGS optimization procedure of the FH molecule.
@@
Figure 3: Estimated error for the value of the force exerted on a) a hydrogen atom (labeled H1) along the O-H1 bond of the water H2O molecule and b) a boron ion (labeled B1) along the z-direction of the $\omega$-BN crystal. The evolution is monitored with respect to changes in their position, using different approximations.
@@
In panel a), the total energy is found to decrease during the relaxation, for every method considered. After only two to three BFGS steps, the variation in $\text{E}_{\text{tot}}$ becomes minimal. This evolution correlates well with the total force, which also reduces over the relaxation steps in panel b). While the $\text{F}_{\text{tot}}$ indicator is initially large, its value decreases significantly for RPA methods after only two BFGS iterations. The decay is more continuous for EXX and HF[PBE] but agrees well with the HF calculations. In the QE code, the default convergence threshold for the geometry is set to 0.0001 Ry on the total energy, and 0.001 Ry/a0 on each of the 3MM individual forces. In the situation considered here, the threshold on the force is reached at the 4th iteration for the RPA-based methods, and at the 5th BFGS step for the HF-based methods. This variation agrees well with the bond distance changes visible in panel c). The F-H bond length is found converged within 0.0005 Å from the 4th relaxation step for RPA[PBE] and RPA, and the 5th iteration for the exchange-only methods. This analysis confirms that our scf and nscf force implementation allows for a smooth optimization of the geometry using the BFGS algorithm.

### v2

@@ I Introduction
The analytical forces in the RPA have been considered in a few works. In Refs. [24, 25, 26, 27, 28, 29, 30], the analytical non-self-consistent RPA forces (i.e. the analytical gradient of the RPA energy evaluated with orbitals and eigenvalues from a semi-local approximation) were implemented in the context of Gaussian basis sets, and tested on a variety of molecular systems. Recently, Thierbach et al. [31, 32] used Gaussian basis sets to evaluate RPA forces at self-consistency, solving the optimized effective potential (OEP) equation [33, 34, 35, 36, 37, 38, 39, 40]. While Gaussian functions are efficient for describing finite-size systems such as molecules, the use of plane waves and pseudopotentials is often more convenient in solids. To date, only one work has reported non-self-consistent RPA forces with plane waves in the projector-augmented-wave framework [41].
In this work, we derive analytical expressions for the self-consistent and non-self-consistent evaluation of RPA forces, in the presence of nonlocal external potentials, such as those arising from norm-conserving pseudopotentials [42]. The force expressions are implemented in the context of plane-wave basis sets within the ACFDT package [43, 37, 12, 44, 14, 18, 39] of the Quantum ESPRESSO (QE) distribution [45, 46], which exploits an eigenvalue decomposition of the linear density response function for evaluating the RPA energy and potential. We demonstrate that our implementation achieves a high numerical quality, enabling the exploitation of RPA forces in a wide range of applications. We also determine forces within RPAx using finite-difference total energy calculations. The accuracy of the different methods is assessed by comparing structural and vibrational properties of molecules and solids to published results obtained with coupled-cluster single, double, and perturbative triple excitations (CCSD(T)), Diffusion Monte Carlo (DMC), and experiments.
The paper is organized as follows. In Sec. II, we recall the ACFDT total energy expression. In Sec. III, we derive the RPA atomic forces in detail and discuss their implementation. In Sec. IV, we demonstrate the numerical accuracy of the calculated forces. In Sec. V, we present the performance of the RPA and RPAx methods in predicting structural and vibrational properties of a set of molecules (H2, LiH, HF, N2, H2O, NH3, and CH4) and solids (C, Si, and Ge). Finally, we present our conclusions in Sec. VI.
@@ II ACFDT formalism
Within the Born-Oppenheimer (BO) approximation, the potential energy surface is given by the sum of the following contributions: the internuclear interaction energy, $E_{\text{NN}}$, the electronic kinetic energy, $T_{e}$, the electron-nuclei (external) interaction energy, $E_{\text{ext}}$, and the electron-electron interaction energy, $W$. Within the KS-DFT framework, the sum of $W$ and $T_{e}$ is re-expressed as a sum of the kinetic energy of independent KS electrons, $T^{\text{KS}}_{e}$, the Hartree energy, $E_{\text{H}}$, the exchange energy, $E_{\text{x}}$, and the correlation energy, $E_{\text{c}}$. Typically, as within the LDA and GGA, both $E_{\text{x}}$ and $E_{\text{c}}$ are approximated.
Methods such as the RPA and RPAx have an exact description of the exchange energy, while the correlation energy is approximated starting from the formally exact ACFDT formula [6, 7],
$E_{\text{c}}^{\text{ACFDT}}=\int\limits_{0}^{1}d\lambda\int\limits_{0}^{\infty}\frac{du}{2\pi}\,\text{Tr}\bigl\{$ $v\bigl[\chi_{\lambda}(iu)-\chi_{\text{0}}(iu)\bigr]\bigr\}.$ (1)
Here, $v$ is the bare Coulomb interaction and $\lambda$ is a parameter connecting the fictitious KS system of independent particles ($\lambda\!=\!0$) to the physical system ($\lambda\!=\!1$), while keeping the charge density fixed at $\lambda=1$.
The key ingredient in the ACFDT formula is the linear density response function, evaluated for the $\lambda$-dependent system and for the KS system. The latter is defined by the one-particle KS potential
@@
The two response functions, $\chi_{\lambda}$ and $\chi_{\text{0}}$, are connected through the Dyson equation of time-dependent density functional theory [47],
@@
The Hartree and exact-exchange kernels are both linear in $\lambda$ ($f^{\lambda}_{\text{x}}=\lambda f_{\text{x}}$). It is, therefore, possible to integrate over $\lambda$ in Eq. (1) analytically. Combining Eqs. (1) and (3) gives the following expression for the correlation energy in RPAx [10] (or RPA, by setting $f_{\text{x}}=0$)
@@
Note here that $f_{\text{x}}$ is frequency-dependent [49]. In Refs. [12, 14], the use of alternative partial resummations of the RPAx was introduced. Using the same $f_{\text{x}}$-kernel, we can define RPAx(1) from
@@
which corresponds to a truncation of the RPAx(1) series that omits exchange in the screened interaction. Once the $f_{\text{x}}$-kernel is calculated, both variants are straightforwardly obtained. They usually give very similar results to RPAx but, in certain critical cases, their use is preferred. In addition to RPAx(1) and AC-SOSEX, other resummations, such as the AXK (approximate exchange kernel) have been found to be accurate [50, 51].
The response functions involve summations over excitation functions, i.e., products of occupied and unoccupied KS states. In Ref. [43], it was shown that the use of unoccupied states can be completely avoided for the implementation of the RPA correlation energy by solving the following eigenvalue problem
@@
Determining the RPA and RPAx correlation energy requires a starting point. Such non-self-consistent calculations always introduce a bias, which can only be completely suppressed by identifying the optimal KS starting point. This is accomplished through the OEP approach [33, 34, 35, 36, 37, 38, 39, 40], which allows orbital-dependent functionals to work within the standard Kohn-Sham scheme of DFT [54]. It applies not only to RPA but also to any method that includes exact exchange, such as hybrid functionals. For a given xc functional, the OEP equation reads
@@
Both the left- and right-hand sides of Eq. (9) can be written in terms of density responses. This feature was exploited in Ref. [37] to solve the OEP equation, combining DFPT with an iterative algorithm for extracting the xc potential. The RPA correlation term contributing to the left-hand side of Eq. (9) can be evaluated as
@@
The variation of each eigenvalue with respect to $v_{\text{eff}}$ can be further split into different terms, which can all be evaluated using DFPT [37, 39].
Following the eigenvalue problem in Eq. (7), the total energy in RPA/RPAx and the RPA potential have all been implemented within the ACFDT code [43, 37, 12, 44, 14, 18, 39] of the QE distribution [45, 46]. Building on these capabilities, we have, in this work, implemented the force expression, as described in the next section.
@@ III Atomic forces in RPA
Having established a way to calculate the total energy, we can now look at its first-order derivatives with respect to the atomic positions. This yields a set of forces, which are important for studying the stability and excitations of a lattice or molecular structure. The force, $\mathbf{F}_{I}$, exerted on a given atom, $I$, corresponds to the derivative of the BO potential energy surface with respect to the atomic position, $\mathbf{R}_{I}$. Thus, from Sec. II, we have
$\mathbf{F}_{I}=-\frac{\partial E_{\text{NN}}}{\partial\mathbf{R}_{I}}-\frac{\partial T^{\text{KS}}_{e}}{\partial\mathbf{R}_{I}}-\frac{\partial E_{\text{ext}}}{\partial\mathbf{R}_{I}}-\frac{\partial E_{\text{H}}}{\partial\mathbf{R}_{I}}-\frac{\partial E_{\text{xc}}}{\partial\mathbf{R}_{I}}.$ (11)
@@
is the external (nuclear) potential. The first term in Eq. (12) corresponds to the Ewald contribution, which is linked to the internuclear repulsion [55]. The second term describes the interaction between the nuclei and the electronic density. The third and fourth terms are the only ones that depend directly on the xc functional. In the next two subsections, we will adapt Eq. (12) to RPA at self-consistency and to non-self-consistent RPA, starting from a local or semi-local approximation.
@@ III.1 Forces at self-consistency
As mentioned previously, in order to avoid starting-point biases, a self-consistent (scf) solution should ideally be considered. Furthermore, self-consistency simplifies the expression for the force given in Eq. (12), thanks to the validity of the Hellmann-Feynman theorem (HFT) [56, 57]. At self-consistency, the xc potential is equal to the functional derivative of the xc energy with respect to the charge density. By applying the chain rule on the fourth term of Eq. (12) and using the fact that, at self-consistency, the density and xc potential are identical in both the third and fourth term, it is easy to see that these two contributions cancel. We can thus write
@@
This is the textbook Hellmann-Feynman force formula, used to compute atomic forces at the LDA and GGA levels of theory.
@@
The effective KS potential, $v_{\text{eff}}(\mathbf{r}^{\prime},\mathbf{r}^{\prime\prime})$, is now allowed to have a (static) nonlocal component coming from Eq. (15). Focusing on the RPA functional, the fourth term of Eq. (12) becomes
@@
The collection of eigenvalues $\{a_{\alpha}\}$ and their variations are obtained directly from the calculation of the RPA correlation energy and potential. The detailed expression for $\delta a_{\alpha}(iu)/\delta v_{\text{eff}}$ when $v_{\text{eff}}$ is local is given in Refs. [37, 39]. For a nonlocal $v_{\text{eff}}$, as in the present case, off-diagonal components also need to be calculated. Fortunately, this does not cause any additional complication. As a result, implementing the calculation of $\partial a_{\alpha}(iu)/\partial\mathbf{R}_{I}$ in the existing scf implementation is relatively straightforward. The computation of the extra scf-RPA correlation force term is also convenient as it can be done together with the computation of the RPA correlation potential. Furthermore, since only the pseudopotential is involved in Eq. (18), the density and eigenvalue variations correspond to bare responses, which facilitates their evaluation.
@@ III.2 Non-self-consistent forces
The force expression in Eq. (12) is valid for any KS starting point based on a local xc potential. For ease of discussion, we will hereafter focus on a particular starting point, the PBE approximation [60, 61]. In this case, the second term in Eq. (12) is evaluated with the PBE density, and the third force term is evaluated with the PBE xc potential. The derivative of $n^{\text{PBE}}$ with respect to $\mathbf{R}_{I}$ is determined from the first-order variation of the occupied PBE orbitals, which can be computed from the self-consistent Sternheimer equation,
@@
where $\hat{H}_{\rm KS}$ stands for the PBE KS Hamiltonian. For a given occupied band $\nu$, the quantity $\varepsilon_{\nu}$ corresponds to its KS eigenvalue, while $\Delta\phi_{\nu}$ is its first-order response to the applied perturbation, $\Delta\mathbf{R}_{I}$. The quantity $\Delta\hat{V}$ stands for the response of the KS potential to the same perturbation. We note that, since self-consistency is required, the only feasible starting points are those from LDA and GGA functionals, for which the DFPT formalism is implemented within QE [53].
@@
Figure 1: Change in the total force for the fluorane molecule according to a) HF, HF[PBE], EXX, and the uncorrected EXX (${\bf F}^{\rm scf\text{-}EXX}_{\rm x}$), b) RPA, RPA[PBE], and the uncorrected RPA (${\bf F}^{\rm scf\text{-}RPA}_{\rm xc}$).
@@ IV Implementation and testing
In order to judge whether the quality of the forces is sufficient for relaxing geometries using, for example, the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm [65, 66, 67, 68], we have considered the fluorane molecule, HF. The hydrogen atom has no core electron and local PPs are found to converge well. On the other hand, fluoride has two core electrons and, as such, benefit from using nonlocal PPs.
@@
$\text{F}_{\text{tot}}=\sqrt{\sum_{I=1}^{M}\sum_{\beta=1}^{3}\left(F_{I}^{\beta}\right)^{2}},$ (23)
@@
where $M$ is the total number of atoms and $\beta$ spans the three spatial Cartesian coordinates. The total force takes into account the individual forces exerted on both the fluorine and hydrogen atoms, and should tend toward zero as the optimal geometry is approached. The change in the total force, calculated at different values of the H-F bond distance, is given in Fig. 1, for different approximations. Panel a) focuses on the methods based on Hartree-Fock. The notation ’HF’, also used to designate the fluorane molecule, refers to a self-consistent calculation done using the nonlocal Hartree-Fock potential, as implemented within the PW program of QE. The ’HF[PBE]’ method corresponds to its non-self-consistent version that we have implemented within the ACFDT package, and ’EXX’ refers to the self-consistent local version of HF that solves the OEP equation. Finally, ’EXX $\{-\mathbf{F}_{\text{x}}^{\text{scf-EXX}}\}$’ corresponds to an EXX calculation that omits the extra exchange force contribution derived in Ref. [42] (exchange part of Eq. (18)). For all these calculations, we used a kinetic energy cutoff of 100 Ry and a simulation cell of size 20 bohrs. For the HF and EXX calculations, a tight self-consistency threshold was set in order to ensure a good convergence of the results. The same was done for obtaining the self-consistent solution to the Sternheimer equation, necessary for computing the forces at the HF[PBE] level (see Eq. (20)).
@@
The HF molecule is found to have an equilibrium bond distance of 0.8924 Å using HF[PBE], 0.8948 Å for EXX, and 0.8952 Å for HF. The minimum in $\text{F}_{\text{tot}}$ is found, for all three methods, at the same position as the minimum in total energy, as expected. However, when we omit the extra exchange force term due to the nonlocal PP, the optimal bond length is about 0.03 Å larger than expected, with an overall error in the total force of about 0.085 Ry/a0. When inspecting the individual forces exerted on each atom along a given Cartesian direction, the weight of the extra PP term is found, as expected, larger for the fluorine atom than the hydrogen, 0.11 Ry/a0 against $\times1$0-4 Ry/a0, respectively.
@@
Panel b) in Fig. 1 focuses on RPA using the same computational setup as for the EXX calculations. For the evaluation of the correlation terms, the frequency integration was performed on a grid of 8 imaginary frequency points, and 10 eigenvalues per valence electron ($N_{\rm eig}$=80) of the KS response function were determined (see Eq. (7)). The bond distance is found at 0.9223 Å with RPA[PBE] and 0.9224 Å with RPA. In both cases, the minimum in the total force matches the minimum associated with the total energy, as expected. When removing the extra xc force term (RPA $\{-\mathbf{F}_{\text{xc}}^{\text{scf-RPA}}\}$), the bias in forces leads to a shorter H-F bond. The error in the evaluation is about 0.01 Å, which corresponds to a shift of 0.027 Ry/a0 in the total force. Again, the extra xc force term is found to be the largest for F, about 0.04 Ry/a0, while it is negligible for H.
@@
Given the good agreement found at equilibrium geometry using either the forces or the total energy, we then investigated how well our calculated forces are exploited by the BFGS algorithm to relax the geometry of the fluorane molecule. For this test, the starting H-F bond distance was set to 1.2 Å. During the relaxation procedure, we fixed the position of the hydrogen atom and only let the position of F to vary. The change over the relaxation steps in the total energy, the total force, and the H-F bond distance, are presented in panels a), b), and c) of Fig. 2, respectively.
@@
Figure 2: Evolution of a) the total energy, b) the total force, and c) the bond distance during the BFGS optimization procedure applied on the HF molecule.
@@
Figure 3: Estimated error for the value of the force exerted on a) a hydrogen atom (labeled H1) along the O-H1 bond of the water H2O molecule and b) a boron ion (labeled B1) along the z-direction ($\mathbf{c}$ crystal axis) of the $\omega$-BN crystal. The evolution is monitored with respect to changes in their position, using different approximations.
@@
In panel a), the total energy is found to decrease during the relaxation, for every method considered. After only two to three BFGS steps, the variation in $\text{E}_{\text{tot}}$ becomes minimal. This evolution correlates well with the total force, which also reduces over the relaxation steps in panel b). While the $\text{F}_{\text{tot}}$ indicator is initially large, its value decreases significantly for RPA methods after only two BFGS iterations. The decay is more continuous for EXX and HF[PBE] but agrees well with the HF calculations. In the QE code, the default convergence threshold for the geometry is set to 0.0001 Ry on the total energy, and 0.001 Ry/a0 on each of the 3MM individual forces. In the situation considered here, the threshold on the force is reached at the 4th iteration for the RPA-based methods, and at the 5th BFGS step for the HF-based methods. This variation agrees well with the bond distance changes visible in panel c). The H-F bond length is found converged within 0.0005 Å from the 4th relaxation step for RPA[PBE] and RPA, and the 5th iteration for the exchange-only methods. This analysis confirms that our scf and nscf force implementation allows for a smooth optimization of the geometry using the BFGS algorithm.